# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v1

@@ 3.3.2 LoRA Fine-Tuning Technique
We illustrate LoRA in Figure 7. We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$. We use $d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices. We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
@@
Figure 7: Illustration of LoRA.
@@ 3.3.3 Implementation Details
We use exactly the same training and test sets as in Section 3.2, but the inputs here are the original headlines with the text information. To help Llama-3-8b understand the boundary between two headlines, we concatenate them using the default end of sequence (EOS) token, <|begin_of_text|>, placed between them. For example, for the headline pair The First Headline and The Second Headline, the combined sequence would be, <|begin_of_text|>The First Headline<|begin_of_text|>The Second Headline.
We add a classification head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure 8. Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence. Thus, the last token in the sequence, being informed by all prior tokens, is used for classification. This embedding vector of the last token, output by the last Transformer layer and having a dimension of 4096, is processed through a linear layer producing an output of two dimensions representing the logits. Then the logits are transformed into probabilities by the softmax function, indicating the likelihood of each headline being more impactful. The softmax function is expressed as:
$\text{Softmax}(Z)_{i}=\frac{e^{z_{i}}}{\sum_{j}e^{z_{j}}}$
where $Z$ is the vector of logits, and $i$ indexes the specific class. During fine-tuning, our objective is to minimize the cross-entropy loss on the training set.
@@
We implement LoRA Fine-tuning using the PEFT library developed by Hugging Face. Our LoRA configuration settings include setting the dimension of the low-rank matrices, denoted as $r$, to 4. The lora_alpha parameter, acting as a scaling factor for the low-rank matrices, is set to 4. This parameter controls how much the learned low rank matrix would be weighted. Adjusting lora_alpha allows us to control the adaptation strength, enabling the model to integrate new information effectively without losing its pre-trained knowledge.
Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora). Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$. We set the dimension of the low-rank matrices $r$ to 4. To mitigate the overfitting, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices. Furthermore, we enable 16-bit mixed precision training. Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly. This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.
With this LoRA configuration on Llama-3-8b, the total number of trainable parameters in the model is 1,703,936, which takes only 0.023% of the total 7,506,636,800 parameters. We fine-tune the model for 10 epochs, meaning each sample (pair of headlines) in the training set is used 10 times in the training. A linearly decreasing learning rate scheduler is employed, starting from an initial rate of 2e-5 and decreasing linearly to zero by the end of the fine-tuning process. This strategy helps stabilize the training by gradually reducing the rate of updates to the weights, thus smoothing the convergence and reducing the risk of overshooting minima in the loss landscape.
Lastly, we want to briefly discuss the GPU requirement and monetary cost of implementing LoRA. The GPU requirements for fine-tuning depend on factors such as model size, batch size, optimization algorithm, and floating-point precision. In our experiment, without LoRA, full-parameter fine-tuning would require over 100 GB of GPU memory. However, LoRA fine-tuning only requires 30 GB of GPU memory. To provide a more straightforward demonstration of the benefit of the saved GPU requirement, we showcase two GPU models that are commonly used for LLM training. Without LoRA, the NVIDIA H100 GPU with nearly the largest GPU memory, 80 GB is insufficient for fine-tuning. In contrast, with LoRA, a more affordable NVIDIA V100 GPU with 32 GB of memory is sufficient. The price of an NVIDIA H100 GPU is ten times higher than an NVIDIA V100 GPU in May 2024. Thus, using LoRA significantly reduces the cost, making fine-tuning more accessible to researchers and practitioners.
@@ 3.3.4 Performance Analysis of Fine-tuning-based Approaches
We report the loss curve and accuracy curve during the LoRA fine-tuning process in Figure 9. We observe a slightly better performance of 82.79% accuracy compared to the best result (82.33% accuracy) using OpenAI embeddings in Section 3.2. It is important to note that, if compared under the same LLM, LoRA fine-tuned Llama-3-8b outperforms the logistic classifier using embeddings from Llama-3-8b (around 77% accuracy).
@@
Figure 9: Loss curve and accuracy curve as the number of training epochs increases. We repeat the fine-tuning process three times and display the average loss and accuracy. For each repetition, we consistently used the same training and test sets, ensuring that any randomness in the results came solely from the fine-tuning process, not from variations in the training and test set splits. The final accuracies of the three replicates on the test set are 83.04%, 82.73%, and 82.59% respectively. We add the best performance (dash line) obtained using OpenAI’s embedding in Section 3.2 for a comparison.
From Figure 9, we observe that the loss evaluated on the training set is consistently lower than the loss on the test set, and the accuracy on the training set is consistently higher than the accuracy on the test set, indicating overfitting. Overfitting refers to the phenomenon that occurs when a machine learning model learns the details and noise in the training data to the extent that it negatively impacts the model’s performance on new, unseen data. This often happens because the model becomes too complex or because the training set is too small, causing the model to memorize specific examples and noise rather than general patterns. As a result, the model performs exceptionally well on the training data but fails to generalize, leading to relatively poor performance on test data.
@@
In our case, the Llama-3-8b model is inherently complex. Although using LoRA reduces the number of trainable parameters to 0.023% of the total parameters, the model remains complex. Common approaches to address overfitting include reducing model complexity, adding regularization, and increasing the dataset size. We experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay, in an attempt to mitigate overfitting. However, these modifications generally reduce performance on the training set without yielding significant improvements on the test set. We suspect that overfitting could be further alleviated, and test performance improved, with a larger and more diverse dataset. The insight for practitioners is that continuously updating the model with new data, such as through A/B testing on a news website to collect more pairwise comparison data, may lead to better generalization and performance over time. Instead of fine-tuning a model version and continuously using it, we suggest that companies keep updating the model as they collect larger datasets.
After fine-tuning the model, several cost-efficient techniques are available for deploying it for inference, significantly reducing computational costs and hardware requirements. For instance, llama.cpp allows LLM models to be deployed on machines without a GPU, making implementation more affordable. Moreover, model compression techniques such as quantization, sparsification, and knowledge distillation can drastically cut inference computation costs while maintaining, or only slightly affecting, model performance. Our approach ensures that the fine-tuned model is not only cost-efficient but also robust and secure when hosted on our own servers. For a comprehensive survey on efficient LLM deployment and inference techniques, please refer to (zhou2024survey).
@@ 4 Our Approach: LLM-Assisted Online Learning Algorithm
So far, we have seen that pure-LLM-based approaches can help with headline selection with reasonably high accuracy (around 82% for the embedding-based and LoRA fine-tuning approaches). The main advantage of these pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is better and go with it. Thus, there is no need to use any traffic for experimentation. The downside, of course, is that these models are not fully accurate—in around 18% of the cases, they are likely to mispredict which headline is better. This can of course, lead to lower clicks and revenues for the firm. [^22] It should be noted that even this accuracy is calculated under the assumption that the results from the A/B tests are the ground truth. However, even if we were to reach 100% accuracy on this dataset, we still may not reach 100% accuracy in practice because of Type I errors in the creation of significant samples. Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
@@
On the other hand, we have the current status quo, which is the use of experimentation-based approaches to choose headlines/content. As mentioned in Section 1, there are two broad experimentation-based methods—(1) A/B tests and (2) Online learning algorithms or bandits. In particular, the standard practice at Upworthy during the time of the data collection was A/B tests, where the firm allocated a fixed amount to different headlines and then chose the best-performing one for the remaining traffic. This approach is also known as Explore and Commit (E&C) and is commonly used in the industry due to its simplicity and low engineering cost; see chapter six of (lattimore2020bandit). The main benefit of this approach is that, with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased; as such it is the gold-standard approach for identifying the best arm (or headline in this case). The downside, of course, is that by assigning traffic to all the headlines during the exploration phase or A/B test, the firm incurs high regret (i.e., low profits since some of headlines in the test are likely have low CTRs). Therefore, many modern technology and media firms such as NYTimes and Yahoo use online learning algorithms or bandits that leverage the explore-exploit paradigm to lower the experimentation cost while still learning (lattimore2020bandit). These algorithms move traffic away from poorly performing arms/headlines dynamically in each period, based on the observed performance of each arm till that period. Because of this greedy behavior, they tend to switch to the best-performing arms quickly, and incur lower regret. Nevertheless, one drawback of these methods is that they start with the assumption that all headlines are equally good, i.e., they suffer from a cold-start problem, which can result in wasting traffic on the weaker headlines early on.
@@
In this section, we present a framework that combines the benefits of LLM-based models with experimentation-based methods. Our key idea is that firms can address the cold-start problem in content experiments by leveraging predictions from LLM-based approaches as bayesian priors in a first step and combine them with online learning algorithm in the second step. Thus, our approach, termed LLM-Assisted Online Learning Algorithm (LOLA), continuously optimizes traffic allocation based on LLM predictions together with the data collected from the online experiment in real-time.
@@
The rest of this section is organized as follows. In Section 4.1, we discuss the general LOLA framework, where we formalize the media firm’s problem and then present our algorithm. Next, in Section 4.2, we discuss the implementation of our proposed algorithm, benchmarks, and the experimental setup. Finally, we present our numerical results in Section 4.3.
4.1 Bandit Framework and LOLA
@@
In the following, we first present the Multi-Armed Bandits (MAB) framework to formalize our problem and then discuss our proposed approach. MAB is a class of sequential decision-making problems where a learner must choose between multiple arms over periods to maximize cumulative reward. In our setting, different arms refer to different headlines associated with the same content, and periods refer to impressions where different headlines with the same content are displayed. Let $[K]:=\{1,2,\dots,K\}$ represent the set of $K$ arms, and $[T]:=\{1,2,\dots,T\}$ represent the entire decision horizon for a test. At each time step $t$, the learner selects an arm $a_{t}\in[K]$ and receives a reward $r_{a_{t}}$ drawn from a probability distribution specific to the chosen arm $a_{t}$. In our setting, we assume each arm has a constant click-through rate (CTR), and the reward is the click feedback, which follows the Bernoulli distribution with the probability equal to the CTR [^23] This reflects a stochastic bandit setting rather than a Bayesian bandit setting. However, it is important to note that our proposed algorithms can also be applied in a Bayesian setting by changing the definition of regret and the process of generating reward feedback. For completeness, we also present a Thompson sampling version of our proposed LOLA algorithm in Web Appendix C..
@@
We first introduce the most common framework, regret minimization. The goal in this framework is to minimize cumulative regret, which is defined as the difference between the reward obtained by always pulling the optimal arm and the reward actually accumulated by the learner. Mathematically, the regret $R_{T}$ after $T$ trials is given by:
$R_{T}=\max_{k\in[K]}\sum_{t=1}^{T}r_{a^{*}}-\sum_{t=1}^{T}r_{a_{t}},$
@@
As a complement to regret minimization, we also tackle the problem using the Best Arm Identification (BAI) framework. The goal of BAI is to identify the arm with the highest reward within a given limited number of pulls. Formally, given a fixed budget $B$, we aim to find the arm $a^{*}$ that maximizes the expected reward after $B$ trials.
@@
Before we introduce our LLM-assisted online learning algorithm, we introduce the widely applied pure learning algorithm without LLMs, the Upper Confidence Bound (UCB) algorithm. The UCB stands out because it is provably asymptotically optimal. The UCB policy is straightforward and selects the arm that maximizes upper confidence bound on the estimated rewards, specifically:
$a_{t}=\arg\max_{k\in[K]}\left(\bar{\mu}^{t}_{k}+\sqrt{\frac{2\log(1/\delta)}{n^{t}_{k}}}\right),$
@@
where $\bar{\mu}^{t}_{k}$ is the estimated mean reward of arm $k$ up to time $t$, $n^{t}_{k}$ is the number of times arm $k$ has been played up to time $t$, and $\delta$ is called the confidence level, which quantifies the degree of certainty. The intuition behind UCB’s effectiveness in balancing exploration and exploitation lies in its construction. By adding a confidence term $\sqrt{{2\log(1/\delta)}/{n^{t}_{k}}}$ to the estimated mean reward $\bar{\mu}_{k}$, UCB ensures that arms with fewer selections (higher uncertainty) are given a chance to be explored. As time progresses and more data is collected, the confidence term diminishes, allowing the algorithm to exploit arms with higher observed rewards. This dynamic adjustment enables the UCB algorithm to systematically explore under-explored arms while exploiting arms with high rewards, thus effectively addressing the exploration-exploitation trade-off inherent in the MAB problem. However, we need to set the key hyperparameter $\delta$ such that it should be small enough to ensure optimality with high probability, but not too small that suboptimal arms are over-explored. Typically, one can set the confidence term as $\alpha\sqrt{\log t/n^{t}_{k}}$ and fine-tune the scale term $\alpha$ for specific problems.

### v2

@@ 3.3.1 LoRA Fine-Tuning Technique
We illustrate LoRA in Figure 4. We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$. Specifically, $W_{Q}$ is the weight matrix for the query vectors, $W_{K}$ is for the key vectors, and $W_{V}$ is for the value vectors in the attention mechanism of the Transformer architecture. These matrices store most of the information and knowledge learned by LLM, and are crucial for calculating the attention scores and subsequently determining the relevance of different tokens in the input sequence. For more technical background, we refer to Web Appendix §$\S$§D.1. We use $d\times d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices. We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
@@
Figure 4: Illustration of LoRA Fine-Tuning.
@@ 3.3.2 Implementation Details
We use exactly the same training and test sets as in §$\S$3.2, but the inputs here are the original headlines with the text information. To help Llama-3-8b understand the boundary between two headlines, we concatenate them using the default end of sequence (EOS) token, <|begin_of_text|>, placed between them. For example, for the headline pair The First Headline and The Second Headline, the combined sequence would be, <|begin_of_text|>The First Headline<|begin_of_text|>The Second Headline.
We add a classification head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure 5. Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence. Thus, the last token in the sequence, being informed by all prior tokens, is used for classification. This embedding vector of the last token (which is 4096-dimensional), output by the last Transformer layer, is processed through a linear layer producing a two-dimensional logit vector, denoted by $Z=(Z_{1},Z_{2})$. The $Z$ vector is then transformed into probabilities by the softmax function, indicating the likelihood of each headline being catchier, i.e.,
$\mathbb{P}\{\text{i-th headline is catchier}\}=\frac{e^{Z_{i}}}{e^{Z_{1}}+e^{Z_{2}}}\mathrm{~{}for~{}}i=1,2.$
@@
During fine-tuning, our objective is to minimize the cross-entropy loss on the training set. We implement LoRA fine-tuning using the PEFT library developed by Hugging Face. We set the dimension of the low-rank matrices as $r=4$. Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora). Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$. To mitigate over-fitting issues, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices. [^14] Furthermore, we enable 16-bit mixed precision training. Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly. This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.
With this LoRA configuration on Llama-3-8b, the total number of trainable parameters in the model is 1,703,936, which translates to only 0.023% of the total 7,506,636,800 parameters. We fine-tune the model for 10 epochs, meaning each sample (pair of headlines) in the training set is used 10 times in the training. A linearly decreasing learning rate scheduler is employed, starting from an initial rate of 2e-5 and decreasing linearly to zero by the end of the fine-tuning process. This strategy helps stabilize the training by gradually reducing the rate of updates to the weights, thus smoothing the convergence and reducing the risk of overshooting minima in the loss landscape.
@@ 3.3.3 Performance Analysis of Fine-tuning-based Approaches
Figure 6: Loss curve and accuracy curve as the number of training epochs increases. We repeat the fine-tuning process three times and display the average loss and accuracy. For each repetition, we consistently used the same training and test sets, ensuring that any randomness in the results came solely from the fine-tuning process, not from variations in the training and test set splits. The final accuracies of the three replicates on the test set are 83.59%, 82.70%, and 83.93% respectively, and 83.40% on average. We add the best performance (red dash line) obtained using OpenAI’s embedding in §$\S$3.2 for comparison.
We report the loss curve and accuracy curve during the LoRA fine-tuning process in Figure 6. We observe an accuracy of 83.40% accuracy on the test data, which is marginally better than the best result (82.50% accuracy) using OpenAI embeddings in §$\S$3.2. However, such marginal improvements should not undermine the utility of fine-tuning. Comparing the fine-tuning of Llama-3-8b with the use of OpenAI embeddings is not straightforward due to inherent differences in models’ capability. To highlight the benefits of fine-tuning, we observe that the MLP classification model using embeddings derived from Llama-3-8b (79.40% accuracy in Table 6) performs worse than LoRA fine-tuned Llama-3-8b (83.40% accuracy). This demonstrates that fine-tuning Llama-3-8b can enhance performance compared to Llama-3-8b embeddings.
@@
Similar to the case of the embedding-based classification model, we notice that the fine-tuning approach also suffers from some overfitting issues; see Figure 6. This is because Llama-3-8b is inherently complex; even though LoRA reduces the number of trainable parameters to 0.023% of the total parameters, the model remains complex. Common approaches to address over-fitting include reducing model complexity, adding regularization, and increasing the dataset size. We experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay in an attempt to mitigate over-fitting. However, these modifications generally reduced performance on the training set without yielding significant improvements on the test set.
We also consider two other versions of the LoRA fine-tuned model to document their performance:
  • $\bullet$
    Larger dataset: One way to improve performance and alleviate overfitting is to fine-tune the model on a larger dataset that also includes data on headline pairs that are not significantly different from each other. In our setting, including insignificant pairs of headlines expands the dataset from 39,158 pairs to 134,941 pairs. However, even with this expanded training data, accuracy on the test set continues to hover around 83%, which indicates that incorporating more training data does not necessarily lead to better performance, at least for the current dataset. We refer interested readers to Web Appendix §$\S$§D.2.1 for details of this analysis and numerical results.
  • $\bullet$
    DistilBERT Fine-tuning: Before the widespread adoption of the term “LLM”, BERT was considered one of the best language models available, and it is still widely used as a benchmark in many NLP tasks. To provide a comparison with these earlier models, we therefore fine-tune DistilBERT – a distilled version of BERT that retains over 95% of BERT’s performance while containing 40% fewer parameters (sanh2019distilbert; devlin2018bert). Fine-tuned DistilBERT provides an accuracy of about 75% on the test dataset, which is good, but still lower than that of the fine-tuned Llama-3-8b model. Please see Web Appendix §$\S$§D.2.2 for details on the implementation and performance of the fine-tuned DistilBERT.
In sum, we find that using more data and/or different base models does not lead to significant improvement in performance in our setting.
3.4 Summary of Performance of Pure-LLM Based Methods
We summarize the performance of various pure-LLM-based methods in Table 7. We also compare the accuracy of human respondents on the same task. Details of the survey questions and the accuracy analysis of human respondents are shown in Web Appendix §$\S$§A.
Method Accuracy Notes
Human respondents 51.58% On 4,571 human labeled pairs
Prompt-based approach $<$65% GPT-4 with in-context learning, evaluated on 1,000 random samples
Classification model with LLM text embeddings 82.50% OpenAI-3072E embedding model with logistic regression
Fine-tuned LLM model 83.40% LoRA fine-tuned Llama-3-8b
Table 7: Summarization of performance of various pure-LLM-based methods and comparison with survey results. All results in this table have been reported before. We only report the best performance of each method.
There are four main takeaways from these comparisons. First, human respondents cannot easily predict which headline would be more engaging and likely to lead to higher CTRs. This is consistent with the fact that media firms (and even expert editors) rely on experimentation to decide which headlines and articles to display to users. Second, while simple prompt-based approaches like in-context learning with examples are marginally better than purely random guesses or human predictions, they nevertheless do not provide sufficiently high accuracy. Third, both embedding-based models and fine-tuning perform equally well (at least in our setting), and provide equally good performance with around 83% accuracy. They each have their pros and cons. The main advantage of the logit classification model with OpenAI embeddings is that it is extremely easy to implement and use in business settings since it does not require any significant model training/deployment. However, embeddings from the OpenAI models are proprietary, and as such, we lack visibility into the exact nature of the embeddings. In contrast, fine-tuning approaches use open-source models such as Llama-3-8b, and we have access to the underlying parameters of the LLM. However, this transparency comes with significant costs – fine-tuning LLMs can be computationally expensive and requires non-trivial training and deployment skills that not all businesses may have or want to invest in. The fourth and most important takeaway is that while pure-LLM-based approaches are informative, they are unable to reach the accuracy of experimental approaches. That is, firms cannot replace content experimentation with LLM-based predictions.
Finally, we discuss the monetary costs associated with three LLM-based approaches. The cost of prompt-based, embedding-based, and LoRA fine-tuning in our analysis are around \$20, \$0.18, \$3, respectively. The cost of prompt-based and embedding-based comes from the access to OpenAI’s API in order to get responses and text embeddings. The cost of fine-tuning comes from renting a cloud computing server equipped with one NVIDIA A100 GPU. Overall, we find that the prompt-based approaches are the most expensive, while embedding and fine-tuning are the cheapest in terms of cost. However, in terms of run time and effort, fine-tuning is more costly. The detailed comparison of the cost, run-time, and engineering effort required for each pure-LLM method are discussed in detail in Web Appendix §$\S$§E.
@@ 4 Our Approach: LLM-Assisted Online Learning Algorithm
So far, we have seen that pure-LLM-based approaches can help with headline selection with reasonably high accuracy (around 83% for the embedding-based and LoRA fine-tuning approaches). The main advantage of pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is better and go with it. Thus, there is no need to use any traffic for experimentation. The downside, of course, is that these models are not fully accurate – in a pairwise comparison, in around 17% of the cases, they are likely to mispredict which headline is better. This can, of course, lead to lower clicks and revenues for the firm. [^15] It should be noted that even this accuracy is calculated under the assumption that the results from the A/B tests are the ground truth. However, even if we were to reach 100% accuracy on this dataset, we still may not reach 100% accuracy in practice because of Type I errors in the creation of significant samples. Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
@@
On the other hand, we have the current status quo, which is the use of experimentation-based approaches to choose headlines/content. As mentioned in §$\S$1, there are two broad experimentation-based methods—(1) A/B tests and (2) Online learning algorithms or bandits. In particular, the standard practice at Upworthy during the time of the data collection was A/B tests, where the firm allocated a fixed amount to different headlines and then chose the best-performing one for the remaining traffic. This approach is also known as Explore and Commit (E&C) and is commonly used in the industry due to its simplicity and low engineering cost; see Chapter 6 of (lattimore2020bandit). The main benefit of this approach is that, with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased; as such, it is the gold-standard approach for identifying the best arm (or headline in this case). The downside, of course, is that by assigning traffic to all the headlines during the exploration phase or A/B test, the firm incurs high regret (i.e., low profits since some headlines in the test are likely to have low CTRs). Therefore, many modern technology and media firms such as NYTimes and Yahoo use online learning algorithms or bandits that leverage the explore-exploit paradigm to lower the experimentation cost while still learning (lattimore2020bandit). These algorithms move traffic away from poorly performing arms/headlines dynamically in each period, based on the observed performance of each arm till that period. Because of this greedy behavior, they tend to switch to the best-performing arms quickly and incur lower regret. Nevertheless, one drawback of these methods is that they start with the assumption that all headlines are equally good, i.e., they suffer from a cold-start problem, which can result in wasting traffic on the weaker headlines early on.
@@
In this section, we present a framework that combines the benefits of LLM-based models with experimentation based methods. Our key idea is that firms can address the cold-start problem in content experiments by leveraging predictions from LLM-based approaches as CTR priors in the first step and combining them with an online learning algorithm in the second step. Thus, our approach, termed LLM-Assisted Online Learning Algorithm (LOLA), continuously optimizes traffic allocation based on LLM predictions together with the data collected from the online experiment in real time.
@@
The rest of this section is organized as follows. In §$\S$4.1, we present the media firm’s problem and the standard bandit framework, and then in §$\S$4.2, we present our LOLA framework. Next, in §$\S$4.3, we discuss the implementation of our proposed algorithm, benchmarks, and the experimental setup. We present our numerical results in §$\S$4.4, and finally, in §$\S$4.5, we show how our LOLA approach can be easily generalized to a wide variety of settings and present a set of natural variants and extensions of the basic LOLA framework.
4.1 Bandit Framework
@@
Multi-Armed-Bandits (MAB) is a class of sequential decision-making problems where a learner must choose between multiple arms over periods to maximize cumulative reward. While there is a large literature on it in the machine learning and statistics community, the literature in marketing is small, but growing; researchers have used it for optimizing website design, pricing, advertising, and to learn consumer preferences (hauser_etal_2009; schwartz2017customer; misra_etal_2019; liberali_ferecatu_2022; aramayo_etal_2023; jain_etal_2024). In our setting, different arms refer to different headlines associated with the same article, and periods refer to impressions where different headlines with the same article are displayed. Let $[K]:=\{1,2,\dots,K\}$ represent the set of $K$ arms, and $[T]:=\{1,2,\dots,T\}$ represent the entire decision horizon for a test. At each time step $t$, the learner selects an arm $a_{t}\in[K]$ and receives a reward $r_{a_{t}}$ drawn from a probability distribution specific to the chosen arm $a_{t}$. We assume each arm has a constant click-through rate (CTR), and the reward is the click feedback, which follows the Bernoulli distribution with a probability equal to the CTR.
@@
The learner’s goal is to minimize cumulative regret, which is defined as the difference between the reward obtained by always choosing the optimal arm and the reward actually accumulated by the learner under the policy played by the learner. Mathematically, the regret $R_{T}$ after $T$ trials is given by:
$R_{T}=\max_{k\in[K]}\sum_{t=1}^{T}r_{a^{*}}-\sum_{t=1}^{T}r_{a_{t}},$ (1)
@@
In MAB settings with finite arms, the standard algorithm that is typically used is the Upper Confidence Bound (UCB) algorithm (auer2002finite). UCB stands out among all online learning algorithms because it is provably asymptotically optimal. The UCB policy is straightforward – it selects the arm that maximizes the upper confidence bound on the estimated rewards. Specifically:
$a_{t}=\arg\max_{k\in[K]}\left(\bar{\mu}^{t}_{k}+\sqrt{\frac{2\log(1/\delta)}{n^{t}_{k}}}\right),$ (2)
@@
where $\bar{\mu}^{t}_{k}$ is the estimated mean reward of arm $k$ up to time $t$, $n^{t}_{k}$ is the number of times arm $k$ has been played up to time $t$, and $\delta$ is the confidence level, which quantifies the degree of certainty. The intuition behind UCB’s effectiveness in balancing exploration and exploitation lies in its construction. By adding a confidence term $\sqrt{{2\log(1/\delta)}/{n^{t}_{k}}}$ to the estimated mean reward $\bar{\mu}_{k}$, UCB ensures that arms with fewer selections (higher uncertainty) are given a chance to be explored. As time progresses and more data is collected, the confidence term diminishes, allowing the algorithm to exploit arms with higher observed rewards. This dynamic adjustment enables the UCB algorithm to explore under-explored arms while exploiting arms with high rewards systematically, thus effectively addressing the exploration-exploitation trade-off inherent in the MAB problem. However, we need to set the key hyperparameter $\delta$ such that it is small to ensure optimality with high probability, but not so small that suboptimal arms are over-explored; see Chapter 7.1 (lattimore2020bandit) for a detailed discussion. Typically, we set the confidence term as $\alpha\sqrt{\log t/n^{t}_{k}}$ and fine-tune the scale term $\alpha$ for specific problems.
4.2 LOLA
We now discuss our approach LOLA, and how the standard UCB approach discussed above can be extended to include information from LLMs in the first step preceding active experimentation. LOLA has two key steps, and we describe each in detail.